OpenStudy (anonymous):

How many 2-digit whole numbers are increased by 18 when their digits are reversed?

OpenStudy (blurbendy):

To represent a 2-digit whole number we can write that as: 10t + u where t represents "tens" and u represents "ones" Then we can add 18 because that's what the question says: (10t+u) + 18 to represent the reverse we just include the opposite of 10t + u, which is 10u + t (10t+u) + 18 = (10u+t) solve: 10t+u+18 = 10u+t 9t-9u+18 = 0 t - u + 2 = 0 u = t + 2 If t = 1, u = 3, then 13 is a solution Can you figure out the rest? If t = 2, u = 4 If t = 3, u = 5 If y = 4, u = 6 If y = 5, u = 7 If y = 6, u = 8 If y = 7, u = 9

OpenStudy (perl):

nice solution :)